OpenStudy (anonymous):

64x^8+x^2 factoring polynomials

OpenStudy (anonymous):

take out x^2, then use sum of cubes.

jimthompson5910 (jim_thompson5910):

Factor out the GCF x^2 to get 64x^8+x^2 = x^2(64x^6+1) Now factor 64x^6+1 64x^6+1 (4x^2)^3+1^3 (4x^2+1)((4x^2)^2-(4x^2)(1)+1^2) (4x^2+1)(16x^4-4x^2+1) So 64x^6+1 factors to (4x^2+1)(16x^4-4x^2+1) ----------------------------------- So 64x^8+x^2 factors to x^2(4x^2+1)(16x^4-4x^2+1) In other words, 64x^8+x^2 = x^2(4x^2+1)(16x^4-4x^2+1)